Find the Domain
f(x)=sqrt(x^2-4)
x^2-4>=0
x^2>=4
Answer x>=2
These are the options to choose from and I'm not sure how to put it in this form. I don't know when to use () or [].
[-2,2]
[0,infinity)
(-2,2)
(-infinity,-2)U(2,infinity)
[-infinity,-2]U[2,infinity]
(0,infinity)
[-infinity,-2)U(2,infinity]
All real numbers
None of these
(-infinity,-2]U[2,infinity)

Answer by stanbon(75887)

Find the Domain
f(x)=sqrt(x^2-4)
x^2-4>=0
x^2>=4
Answer x>=2 OR x<=-2
------
(-oo,-2] U [2,+oo)
======================
Cheers,
Stan H.
======================